OpenStudy (anonymous):

Find the infinte limit f(x)= lim->-1^+ (x^2+2)/(x^2-x-2)

OpenStudy (anonymous):

can you clean up the equation so i understand what you are asking?

OpenStudy (anonymous):

\[\large \lim_{x\to-1^+}\frac{x^2+2}{x^2-x-2}\] Substituting directly won't work, since you get \(\dfrac{3}{0}\). To see what happens at the singularity, take values of \(x\) that get successively closer to \(-1\) from the right side. In other words, evaluate the expression for numbers like \(x=-0.9,-0.09,-0.009\), and so on.